(1 point) The histogram below gives the length of service of members of the Department of Biology at a particular university The classes, in years of service, are 0-49,5-9.9, etc, and the vertical axis represents the number of taculty What percent of the department faculity have less than 25 years of service? Answer Biology Faculty 5 1 1520 Years of Service
Answer #1

Total number of faculty=4 + 5 + 8 + 4 + 10 + 8 + 3 + 8 = 50

Number of faculty having less than 25 years of service =

4 + 5 + 8 + 4 + 10 = 31

Therefore, the percent of the department faculty have less than 25 years of service = (31/50) * 100% = 0.62 * 100% = 62%

Answer: 62%